the sum of squares of three distinct real numbers which form an increasing GP is S^2(common ratio r) if the sum of numbers is tS,then
1.If r=3 then t^2 cannot lie in
A.[1/3,1]    B.(1,2)    C.[1/3,3]    D.(1,3)
2.if t^2=2 then the value of common ratio r is greater than
A. 9     B. 4   C. 2   D. 3

Grade:11

1 Answers

Vikas TU
7 years ago
a,b,c be in G.P.
then,
b/a = c/b = r
and 
condtns. given are:

a^2 + b^2 + c^22 = S^2
and
a + b + c = tS
 
  1. if r = 3
then
b = 3a and c = 9a
therefore the eqn.s becomes:
 
13a = tS 
and 
91a^2 = S^2
dividing them it comes,
t = 13/root91
or
t^2 = 169/91 which is approx.  ====> 1 point something
that satiffies hence the B,C and D options excluding A.